OpenStudy (anonymous):

fanning + medal Solve the following system of equations. Please show your work to receive full credit. x - y = 13 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice.

OpenStudy (anonymous):

First you solve for x. @HelloGoodmorning

OpenStudy (gorv):

x-y=13 y=x-13 putting y into 2x+y=2 2x+x-13=2 3x=2+13 x=15/3 x=5 y=-8

OpenStudy (anonymous):

put x=13+y in eqn 2 2(13+y)+y=2 26+2y+y=2 3y=-24 y=-8 x=13+y implies x=13-8=5 in this method i replaced one variable in one equation with the second one using one equation to transform my other equation into a single variable one which is much easier to solve

OpenStudy (anonymous):

Let @HelloGoodmorning at least try. :)

OpenStudy (anonymous):

Sorry i'm back

OpenStudy (anonymous):

@giligails what does eqn mean equations?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Alright thanks and thank you loveiskey

OpenStudy (anonymous):

@satellite73 can you tell me what method giligails used please I believe it was subistitution